OpenStudy (anonymous):

"Medal For best best answer"--------- Consider the three points: A=(6,8) B=(1,2) C=(5,7) Detrmine the angle between AB and AC

OpenStudy (amistre64):

define AB and AC, then use the property: \[\vec u\cdot \vec v=|\vec u||\vec v|~\cos(\alpha)\]

OpenStudy (anonymous):